OpenStudy (anonymous):

Would there be no solutions? -x+2z=-9 -x-3y-4z=2 -3x-2y+2z=17

OpenStudy (hba):

@Rayyee Did you try to solve it ?

OpenStudy (hba):

How can you say that there will be no solutions ?

OpenStudy (anonymous):

Because when I plugged it in the Calculator ( you know the Matrix method) it said like error meaning there is no solution?

OpenStudy (hba):

Well there will be solutions,What about trying the substitution method ?

OpenStudy (hba):

@Rayyee Or Is your question related to Matrices ?

OpenStudy (anonymous):

OKay no I tryed doing that so I got -21-42z=77 and 21y+42z=33 it would cancel each other out...

OpenStudy (hba):

yeah adding them up will get you y :)

OpenStudy (anonymous):

oops I ment -21y-42z=77 and 21y+42z=33

OpenStudy (anonymous):

the y val and z value cancels O:

OpenStudy (hba):

Well then there is no solution.

OpenStudy (hba):

@Rayyee I think you are done ?

OpenStudy (anonymous):

ut may you double check X) Because maybe I made a mistake but to me I keep re checking and it seem slike I am still right

OpenStudy (hba):

@Rayyee You are Right then i guess.

jimthompson5910 (jim_thompson5910):

I just answered this question (about 25 minutes ago) and the numbers and equations are the exact same. See this link http://openstudy.com/users/jim_thompson5910#/updates/50ae767ee4b09749ccabe850

OpenStudy (hba):

Good to hear you answered this question :D

OpenStudy (anonymous):

OH COOL thanks!

OpenStudy (hba):

As i Said @Rayyee You Are correct.